A management of a factory claims that 70% of all their employees have been vaccinated. If a sample of 15 people are chosen, find probability that more than 13 employees actually get vaccinated.


1
Expert's answer
2021-11-15T18:59:15-0500

Let "X=" the number of employees who have been vaccinated: "X\sim Bin(n, p)."

Given "n=15, p=0.7, q=1-p=1-0.7=0.3."


"P(X>13)=P(X=14)+P(X=15)"

"=\dbinom{15}{14}(0.7)^{14}(0.3)^{15-14}+\dbinom{15}{15}(0.7)^{15}(0.3)^{15-15}"

"=0.030520038278205+0.004747561509943"

"=0.035267599788148"

"\approx0.0352676"


The probability that more than 13 employees actually get vaccinated is "0.0352676."